Use a proof by contradiction to show that there is no rational number r for which r+r+1=0.

(Assume that r=a/h is a root, where a and b are integers and a/b is in lowest terms.obtain an equation involving integers by multiplying by then look at whether a and bare each odd or

even.
1
Expert's answer
2021-09-09T05:00:39-0400

Use a proof by contradiction to show that there is no rational number "r" for which "r^3 + r + 1 = 0."

Proof

Assume that "r = a\/b" is a root, where "a" and "b" are integers and "a\/b" is in lowest terms.

We see that

if "a" is odd, then "b" is even,

if "a" is even, then "b" is odd.

Substitute


"\dfrac{a^3}{b^3}+\dfrac{a}{b}+1=0"

Multiply by "b^3\not=0"


"a^3+ab^2+b^3=0"

If "a" is odd, "b" is even, then

"a^3"is odd, "ab^2" is even, "b^3" is even. Hence the sum "a^3+ab^2+b^3" is odd. But zero is an even number.

We have a contradiction in this case.


If "a" is even, "b" is odd, then

"a^3"is even, "ab^2" is even, "b^3" is odd. Hence the sum "a^3+ab^2+b^3" is odd. But zero is an even number.

We have a contradiction in this case.


Therefore our assumption leads to the contradiction.

Therefore there is no rational number "r" for which "r^3 + r + 1 = 0."